Find an equation for the nth term of the arithmetic sequence.

-15, -6, 3, 12, ...

an = -15 + 9(n + 1)

an = -15 x 9(n - 1)

an = -15 + 9(n + 2)

an = -15 + 9(n - 1)

show work and explanation show all your work please
1
Expert's answer
2014-05-23T02:57:15-0400
As we can see the first term of this sequence is -15.
The step of the arithmetic progression is a_2 - a_1 = -6 - (-15)=9
Then the nth term is a_n = a_0 + (n-1) d = -15 + 9(n-1.)